### v1

# Read Anywhere Pointed: Layout-aware GUI Screen Reading with Tree-of-Lens Grounding

Read Anywhere Pointed:
Layout-aware GUI Screen Reading with Tree-of-Lens Grounding
Yue Fan University of California, Santa Cruz Lei Ding University of California, Santa Cruz Ching-Chen Kuo eBay Inc. Shan Jiang eBay Inc. Yang Zhao eBay Inc. Xinze Guan eBay Inc. Jie Yang Cybever
Yi Zhang
University of California, Santa Cruz
Xin Eric Wang University of California, Santa Cruz
[…]
[leftmargin=5pt,rightmargin=5pt,backgroundcolor=gray!20,innertopmargin=5pt,innerbottommargin=5pt]coloredquote
Read Anywhere Pointed:
Layout-aware GUI Screen Reading with Tree-of-Lens Grounding
Graphical User Interfaces (GUIs) are central to our interaction with digital devices. Recently, growing efforts have been made to build models for various GUI understanding tasks. However, these efforts largely overlook an important GUI-referring task: screen reading based on user-indicated points, which we name the Screen Point-and-Read (ScreenPR) task. This task is predominantly handled by rigid accessible screen reading tools, in great need of new models driven by advancements in Multimodal Large Language Models (MLLMs). In this paper, we propose a Tree-of-Lens (ToL) agent, utilizing a novel ToL grounding mechanism, to address the ScreenPR task. Based on the input point coordinate and the corresponding GUI screenshot, our ToL agent constructs a Hierarchical Layout Tree. Based on the tree, our ToL agent not only comprehends the content of the indicated area but also articulates the layout and spatial relationships between elements. Such layout information is crucial for accurately interpreting information on the screen, distinguishing our ToL agent from other screen reading tools. We also thoroughly evaluate the ToL agent against other baselines on a newly proposed ScreenPR benchmark, which includes GUIs from mobile, web, and operating systems. Last but not least, we test the ToL agent on mobile GUI navigation tasks, demonstrating its utility in identifying incorrect actions along the path of agent execution trajectories. Code and data: screen-point-and-read.github.io.
Figure 1: Our ToL agent describes the region on the screenshot indicated by a user’s point. Distinguished from other screen reading tools, our ToL agent can output layout-aware descriptions for points anywhere on the screen.
1 Introduction
Graphical User Interfaces (GUIs), dominate our digital interactions with visually rich screenshots featuring colors, icons, texts, and spatial layouts. Recognizing that screenshots are more accessible and intuitive for depicting visual cues, there are growing efforts to build AI agents to interpret GUIs visually Shaw et al. (2023); Deng et al. (2023); Cheng et al. (2024); Hong et al. (2023); You et al. (2024). Existing works, however, overlook the point-based screen reading task, where the input is the coordinate of a user’s indicated point on the screen and the corresponding screenshot, and the output is a descriptive interpretation of the screen region pointed. This task is critical for accessible technology, providing valuable assistance to users with visual impairments and we refer to such a task as the Screen Point-and-Read (ScreenPR) task.
Aiming to solve the ScreenPR task, we introduce the Tree-of-Lens (ToL) agent. Taking advantage of the generalizability of the advanced Multimodal Large Language Models (MLLMs), the ToL agent is built for GUIs screenshots across domains with a point anywhere on the screen as input. The output from the ToL agent is natural language descriptions for both the content of the indicated region and the related information about the screen layout, as shown in Figure 1. Especially, we find that it is critical to articulate the layout between elements in the screenshot for users to gain a comprehensive understanding of the interface and avoid ambiguity. For example, with only content description, one cannot distinguish the two identical “Tumbler pack" shown in Figure 1.
Our ToL agent employs a ToL grounding mechanism, constructing a Hierarchical Layout Tree for each input screenshot to depict its underlying structure. The nodes in the tree represent regions of varying scales. We build this tree using an object detection model trained on our newly collected Android Screen Hierarchical Layout (ASHL) dataset, which contains 50k bounding boxes of hierarchical screen regions from Android screenshots. This model enables an automatic extraction of local and global regions, forming the Hierarchical Layout Tree. After constructing the tree, we extract the target path based on the region of interest, producing lenses as visual prompts for the MLLM, simulating a human-like, gradually refined focus.
To rigorously evaluate our ToL agent against other baselines, we introduce the Screen Point-and-Read (ScreenPR) benchmark with the ScreenPR task that requires the model to generate descriptions based on user-indicated points. This benchmark consists of 650 screenshots from web, mobile, and operating system GUIs, manually annotated with 1,500 target points and regions. With the ScreenPR benchmark, we show that our ToL agent achieves the best performance compared with the baselines of other general and GUI-specialized MLLMs, with over 15% and 30% improvements respectively in terms of the accuracy of content and layout descriptions compared with vanilla text-prompting GPT-4o.
In addition, we also test our ToL agent with trajectories from a mobile GUI navigation agent. By applying ToL agent to describe the point of each action taken in the trajectory, we demonstrate its utility in identifying incorrect actions along the execution path. This capability is pivotal for refining the development of mobile agents, as it provides clear, actionable feedback on navigation decisions. This application not only validates the ToL agent’s effectiveness but also opens up possibilities for strengthening the performance of mobile agents with sophisticated, layout-aware GUI understanding capabilities.
In conclusion, the contributions of our paper are as follows:
  • We develop the Tree-of-Lens (ToL) grounding method and build the ToL agent that generates descriptions that effectively combine both content and layout information of the pointed screen regions.
  • We propose the Hierarchical Layout Tree to represent the underlying hierarchical structure of the screenshot. We collect the Android Screen Hierarchical Layout dataset and train a model to build the Hierarchical Layout Tree.
  • We introduce the Screen Point-and-Read benchmark, which specifically challenges the model to produce accurate descriptions that include both content and layout details.
  • We rigorously test the ToL agent on the Screen Point-and-Read benchmark and we demonstrate its ability to accurately identify incorrect actions in mobile navigation trajectories.
Figure 2: Pipeline of the Tree-of-Lens agent. The Hierarchical Layout Tree is first constructed based on detected global and local regions from the input screenshot. Then, a set of hierarchical lenses with various field widths is generated from the selected target path in the tree and sent as visual prompts to GPT-4o to generate the content and layout descriptions.
2 Related Works
MLLMs for GUI Understanding
Recent advancements in Multimodal Large Language Models (MLLMs) have led to arising a series of highly capable generalist models Liu et al. (2024b); Chen et al. (2023); Ye et al. (2023); Dai et al. (2023); Bai et al. (2023); Liu et al. (2024a). Spurred by the strong general capability of MLLMs, recent works have been focused on utilizing and improving MLLMs for GUI tasks. Some focus mainly on a specific GUI task such as GUI navigation Wang et al. (2024); Yang et al. (2023b); Zheng et al. (2024), GUI referring expression comprehension Cheng et al. (2024), while some aim at a set of GUI grounding and referring tasks simultaneously, such as CogAgent Hong et al. (2023), which is Multi-task Fine-tuned on CogVLM Wang et al. (2023), and Ferret-UI You et al. (2024), ensuing Ferret You et al. (2023); Zhang et al. (2024), tailored for GUI domain. In our work, we focus on the Screen Point-and-Read (ScreenPR) task, a referring task in which inputs are point locations on the screen with the corresponding screenshots, and develop a specific benchmark to evaluate the model’s performance in such tasks. Also, inspired by big challenges that MLLMs are facing in understanding images with complex layout information Fan et al. (2024), we propose the Tree-of-Lens (ToL) agent that generates layout-aware description, which distinguishes from all other GUI agents with GUI referring ability.
Chain-of-Thoughts for Vision
The Chain of Thoughts (CoT) Wei et al. (2022) methodology in natural language processing (NLP) has spurred various inspiring ideas in this direction. Creswell et al. (2022) applied it to interpretable logical reasoning, Wang et al. (2022) focused on enhancing the "exploring-selecting" mechanism in reasoning path, and Yao et al. (2024); Besta et al. (2024) brought in more sophisticated patterns for reasoning. To embrace CoT in the vision area, Xi et al. (2023) uses Chain-of-Look prompting for surgical scene understanding to learn rich semantics from surgical endoscopic videos. In comparison, Chen et al. (2024) designated an iterative step-by-step reasoning manner for VQA problem, letting MLLMs "See-Think-Confirm" to generate a better rationale for the decided answer. Recently, Shao et al. (2024) proposed a multi-turn processing pipeline that dynamically focuses on visual inputs and provides interpretable thoughts. Moving along with the direction, we aim to facilitate MLLMs on the GUI referring task by generating a chain of lenses with varying field widths, ranging from fine-grained to coarse views. The chain of lenses demonstrates the hierarchical layout of the GUI, simulating a human-like sequence of attention, improving MLLMs on producing robust and accurate descriptions with the content and layout information regarding the target point input in the accessible screen reading task.
3 Tree-of-Lens Agent
3.1 Task Definition
We aim to develop an agent to interpret designated regions on GUIs indicated by points, which we refer to as the Screen Point-and-Read (ScreenPR) task. This agent needs to process input consisting of a screenshot $S_{i}$ and a point coordinate $P_{i}$ on the screen, where the point represents a region of the user’s interest. Two target outputs are language descriptions $\hat{D}^{c}_{i}$ for the content of the targeted region and $\hat{D}^{l}_{i}$ for the layout information of the screen related to the targeted region.
3.2 Method Overview
Targeting the ScreenPR task, we propose the Tree-of-Lens (ToL) agent equipped with the ToL grounding mechanism, as shown in Figure 2. Since our agent relies on the pure visual modality of the GUI and leverages the strong visual referring capabilities of MLLMs, such as GPT-4o, it can effectively read anywhere pointed on screens from any domain, including but not limited to web, mobile, and operating systems.
The three distinct design goals distinguish our LoT agent from previous models. First, our LoT agent outputs rich, natural language descriptions rather than fixed template descriptions from accessible screen reading tools such as VoiceOver [^1] https://support.apple.com/guide/voiceover. Second, it can handle any point location on the screen as input by dynamically detecting the pointed region, which might be as small as an icon or a broader area, such as a full-screen window. Last but not least, the descriptions generated by our agent encompass both the content and layout perspectives, detailing the targeted region’s function and position within the GUI’s layout.
3.3 Hierarchical Layout Tree Construction
Hierarchical Layout Tree
The layout of the GUI screenshot usually features an inherent hierarchy decided by GUI source codes. We propose the Hierarchical Layout Tree to represent this hierarchical structure. This tree structure corresponds to each screenshot, with each node representing a square region within the GUI screenshot. Instead of using varied depths of tree structures from the underlying source codes of GUIs, for simplicity, the proposed Hierarchical Layout Tree has a fixed 3-layer depth for every GUI screenshot, with nodes in each layer representing regions of varied hierarchies. The top layer node represents the whole screenshot. The middle layer nodes represent global regions, which encompass larger, comprehensive areas that include multiple related elements, such as panels or groups of controls, forming significant parts of the user interface. The leaf nodes represent local regions, referring to smaller, more specific areas, typically denoting individual interactive elements like buttons, text fields, and icons.
Model and tree construction pipeline
In order to extract the Hierarchical Layout Tree that depicts the hierarchical layout of screenshots without relying on such structured text modality, we train a GUI region detection model to detect the local and global regions for each GUI screenshot, then construct the Hierarchical Layout Trees accordingly. The GUI region detection model is fine-tuned on the DINO detection model with ResNet-50 backbone Zhang et al. (2022); Chen et al. (2019). Once the regions are detected, we construct the Hierarchical Layout Tree, where each predicted global region serves as a node connected to one or more leaves representing the predicted local regions. The connection between global and local regions is determined by the highest IoU between the local and global regions, where for a predicted local region $\hat{R_{i}}$, its parent global region is $\hat{R_{j}}$, $j=\textit{ArgMax}_{j}\textit{IoU}(R_{i},R_{j})$.
Android Screen Hierarchical Layout dataset
To train a detection model for the Hierarchical Layout Tree, we build a dataset named the Android Screen Hierarchical Layout (ASHL) dataset, featuring screenshots and bounding boxes labeled as either global or local regions. There are around 52k bounding boxes from 2,180 screenshots, spanning all 18 different Android applications.
[…]
$R_{i+j}=\textit{Max}(R_{i},R_{j}),$
if $\quad\textit{IoU}(R_{i},R_{j})>0.9$
[…]
Notably, we find that the merged connected nodes $R_{i+j}$ often indicate regions having united semantic contents due to their multiple code attributes. Therefore, we further label those merged nodes with more than one leaf as global regions and the corresponding leaves as local regions:
[…]
$\{R_{i+j}$ $|\textit{len}(\textit{Leaf}(R_{i+j}))>1\}$ (2)
[…]
$\{\textit{Leaf}(R_{i+j})$ $|\textit{len}(\textit{Leaf}(R_{i+j}))>1\},$ (3)
where $\textit{Leaf}(R_{i})$ outputs all the leaves in the tree tracing down from the node of region $R_{i}$.
Figure 3: Example of the lenses generated from the Hierarchical Layout Tree based on a point coordinate. Lens 2 can be seen as a zooming-out from Lens 1. ToL agent’s corresponding output is in Appendix A.
3.4 Target Path Selection and Multi-lens Prompting
Based on the constructed Hierarchical Layout Tree, we next select a target path from the tree and generate a series of image prompts with annotations based on the input screenshot $S_{i}$ and point $P_{i}$. The image series represents different "lenses" with varying field widths and becomes the visual prompt for the MLLM, GPT-4o, which we refer to as Multi-lense prompting.
The process first identifies the smallest local region that contains the input point. Then the connected parent global region is extracted from the Hierarchical Layout Tree [^2] In some rare cases, the input point is outside all local regions, we treat the global regions as local regions in this process and the parent of the global region is the full screenshot.. We next curate two lenses in a sequence that visually articulates this hierarchical relationship from a fine-grained to a coarse view: the first lens shows only the global region from the screenshot while marking the local region with a colored bounding box labeled ‘1’, the target point with a semi-transparent red dot; the second lens includes the complete input screenshot and marks the global region with another bounding box labeled ‘2’. Figure 3 shows an example.
Then a Multi-lens Prompting method is adopted, where the curated set of lenses is sent to GPT-4o with a text prompt which includes the text coordinate of the target point. Some specific instructions are also included in the prompt, asking, "Explain where box 1 is in relation to box 2 and where box 2 is located on the overall screen." In this way, we provide clear visual cues that facilitate the GPT-4o to generate detailed descriptions that reflect the GUI’s content and layout. For more details about the prompt used, please refer to Appendix C.
4 Screen Point-and-Read Benchmark
4.1 Overview
We introduce the Screen Point-and-Read (ScreenPR) benchmark to rigorously evaluate our ToL agent on the ScreenPR task, where descriptions are required based on user-indicated points on screenshots. As shown in Table 1, this benchmark covers a diverse domain of GUIs with a total 650 screenshots. Each screenshot is annotated with about 2 target points and their corresponding local regions on the screenshot. Figure 4i illustrates the normalized coordinates of target points $P_{i}$ in the ScreenPR benchmark, demonstrating thorough and comprehensive coverage of every location on the screen. The regions included have various content, and their area ranges from 1% to over 50% of the screen, as shown in Figure 4ii.
i
ii
Figure 4: (i) shows the normalized locations of target points $P_{i}$ in the input screenshot $S_{i}$. (ii) shows distributions of the area of local regions $R_{i}$ in our ScreenPR benchmark.
Domain Screenshots Target Points Local Regions
Web 199 500 500
Mobile 201 500 500
Operating System 250 500 500
Total 650 1,500 1,500
Table 1: Key Statistics of the ScreenPR Benchmark
4.2 Data Collection
We first gather mobile and web screenshots from Screenspot Cheng et al. (2024) and operating system screenshot from the data explorer videos [^3] https://os-world.github.io/explorer.html of OSWorld Xie et al. (2024). Then we randomly generate 3 or 4 candidates of the target point $P_{i}$ for each screenshot. Then, to maximize the variety of the region indicated by points, we hire 3 students to check these candidate points, where they assess all the candidate points on each screenshot and remove those that represent duplicated regions. Then, the remaining points are regarded as the target point $P_{i}$, and the annotators draw the bounding boxes of the corresponding local regions $R_{i}$. After the annotation, scripts were used to spot errors, such as ensuring that local regions $R_{i}$ included $P_{i}$ inside. The annotated data is then manually examined by authors to ensure quality.
4.3 Evaluation
The ScreenPR benchmark evaluates both the accuracy of the generated content description $\hat{D}^{c}$ and the generated layout description $\hat{D}^{l}$. Since there is no ground truth for the generated descriptions, we involve human evaluations and also propose automated cycle consistency evaluations inspired by the two-agent evaluation in Padmakumar et al. (2022); Fan et al. (2023).
Human evaluation
We conduct surveys with human judges to assess the quality of the generated descriptions. Each survey presents a screenshot with an indicated point, accompanied by descriptions from different screen readers. The judges are asked to rate these descriptions based on three criteria: how well they describe the content, how well they articulate the layout, and their overall preference. Each question had four options, ranging from "very well" to "not at all," which we numerically mapped to accuracy scores of 100%, 66%, 33%, and 0%, respectively.
Cycle consistency evaluation
In order to evaluate screen reader models automatically, we propose the cycle consistency evaluation. The outputs $(\hat{D^{c}},\hat{D^{l}})$ from the screen-reader are each used as the input for an auxiliary model tasked with specific validation tasks, with the alternative model’s performance indicating the quality of the screen-reader’s descriptions.
Specifically, to evaluate content accuracy, GPT-4o serves as the auxiliary model, performing a multi-choice selection task. This alternative model is asked to select one from the multi-choice candidates based on the predicted descriptions $\hat{D^{c}}$. The candidates are four screenshots of different GUI regions cropped based on human annotation, with only one targeted region
[…]
$k_{0}=i,k_{1},k_{2},k_{3}\in N\},$ (5)
where $Q$ is the fixed question and $\textit{Crop}(S_{i},R_{i})$ means the screenshot $S_{i}$ cropped at the region $R_{i}$.
For layout description accuracy, we design another multi-choice selection task for the auxiliary model, GPT-4-turbo. This multi-choice selection task requires the auxiliary model to select the positional relationship of two regions indicated by $P_{i}$ and $P_{i^{\prime}})$ based on corresponding generated layout descriptions $\hat{D^{l}}_{i}$ and $\hat{D^{l}}_{i^{\prime}}$. $P_{i^{\prime}}$, the reference point, is manually annotated to represent a region $R_{i^{\prime}}$ that does not overlap with the local region of the target point $R_{i}$ on the same screen $S_{i}$, especially for the evaluation of the layout description accuracy. The multi-choice selection task presented to the auxiliary model is
[…]
where $Q_{\text{gen}}$ is a question template that needs to be filled with the layout descriptions generated, and $C$ are some fixed choices provided such as ‘upper,’ ‘lower,’ ‘left,’ etc. Based on the ground truth location of $R_{i}$ and $R_{i^{\prime}}$, we calculate the accuracy of the choices made by the auxiliary model, which indicates the screen reader’s proficiency in describing the layout.
Avg. Len. Human evaluation (%) Cycle consistency evaluation (%) Language Similarity
Content Layout Overall Content Acc Layout Acc BERTScore ROUGE
LlaVA-NeXT 51.95 21.69 9.59 21.19 25.43 7.20 85.18 16.52
CogAgent 30.69 22.18 13.33 20.62 25.86 8.00 82.31 15.49
GPT-4o 34.44 32.73 28.12 30.88 35.10 21.87 85.76 15.08
ToL Agent 38.37 41.82 33.42 41.63 74.96 39.67 86.73 19.47
Table 2: Main results. We evaluate our ToL agent against three other baselines with the human evaluation and cycle consistency evaluation of our ScreenPR benchmark. Additionally, we compare the generated content descriptions with human verified content descriptions using language similarity scores. The results show that the ToL agent achieves the best performance.
Content
Acc
Layout
Acc
ToL Agent 75.0 39.7
+ w/o multi-lens visual prompt 71.2 (-3.8) 37.7 (-2.0)
+ w/o target point mark 65.6  (-9.4) 37.6 (-2.1)
+ w/o local & global regions marks 35.1 (-29.9) 21.9 (-18.6)
Table 3: Ablation results. We gradually refine the ablation to exclude various processes in our ToL grounding mechanism for ToL agent on the ScreenPR benchmark. The results show the effectiveness of various components.
5 Experiments
Baselines
We adopt several baselines, including generalist MLLMs: the proprietary GPT-4o OpenAI (2024) and the open-source LlaVA-NeXT Liu et al. (2024a). We also test the CogAgent Hong et al. (2023), a MLLM specifically tuned for GUI tasks. The inputs for baselines are the text prompt with the input point coordinate $P_{i}$ and the screenshot $S_{i}$. We keep prompts for baselines in the same style as the ToL agent, expect that we have to simplify it for CogAgent as we find it necessary for CogAgent to follow the instructions. Detailed prompt design is shown in Appendix C.
Training
We divide the ASHL dataset into 90 % training and 10% evaluation sets for the GUI region detection model training process. Training is performed on four NVIDIA A6000 GPUs with a batch size of 8, spanning 90 epochs. The optimized model demonstrated excellent performance, achieving an Average Precision (AP) of 94.1% and an Average Recall (AR) of 95.9%.
5.1 Main Results
We evaluate our ToL agent against three baselines on the ScreenPR benchmark as shown in Table 2. The results are consistent across both human evaluation and automatic cycle consistency evaluation, showing that our ToL agent achieves the best performance in terms of content and layout description accuracy. Notably, our ToL agent demonstrates over a 15% improvement compared to the GPT-4o model. Furthermore, the overall scores from the human evaluation indicate that descriptions generated by our ToL agent are the most favorable. Interestingly, although the results show that the CogAgent performs slightly worse than LlaVA-NeXT in terms of the content and layout description accuracy, the overall scores from human judges slightly prefer LlaVA-NeXT. We believe this preference is due to the LlaVA-NeXT model’s tendency to produce longer descriptions with more information about the screen, creating a more comprehensive feel.
Additionally, we leverage language similarity scores, including BERTScore and ROUGE-L, to compare the generated content descriptions with human-verified descriptions. These descriptions are captions generated using GPT-4V with the direct input of $Crop(S_{i},R_{i})$, and we manually examined and corrected mistakes. We observe that the overall trend from the language similarity evaluation results matches the content evaluation results from human and cycle consistency evaluations. However, the differences in language similarity scores between models are relatively indistinguishable. We believe this is because the captions are more comprehensive, sometimes redundant, whereas the screen reading output is more focused on delivering the key features of the content to help users understand it.
Content Acc Layout Acc
Human 92.8 70.3
ToL Agent 72.5 (-20.3) 41.3 (-29.0)
ToL w/ GT local regions 85.5 (-7.3) 44.3 (-26.0)
ToL w/ GT global regions 88.2 (-4.6) 46.6 (-23.7)
ToL w/ GT global local regions 89.5 (-3.3) 48.2 (-22.1)
Table 4: Bottleneck analysis. Replacing local and global regions with GT annotations mainly improves content description accuracy, indicating rooms of improvements for both the GUI region generation process and the MLLM leveraged.
Figure 5: Pipeline of employing our ToL agent in verifying the actions from a mobile navigation agent.
5.2 Ablation Study
Effectiveness analysis
To assess the effectiveness of the ToL design, we conduct an ablation study where we progressively refine the ablation to exclude key design components and evaluate these ablated versions on our ScreenPR benchmark using the cycle consistency evaluation, as shown in Table 3. First, we modify our ToL agent to exclude the multi-lens visual prompting. This change necessitates annotating bounding boxes for the local region, global region, and the target point on the complete screenshot as a single visual prompt. Then we further remove the red dot mark from the visual prompt, resulting in a design similar to the Set-of-Mark (SoM) prompting method Yang et al. (2023a). As a result, these changes lead in performance degradation in both content and layout description accuracy, indicating the effectiveness of our visual prompting strategies. Finally, we ablate the input visual prompt to remove marks of local and global regions that we extract from our Hierarchical Layout Tree. This drastic modification results in a significant decrease in performance, underscoring the importance of our Hierarchical Layout Tree in the of the ToL design.
Bottleneck analysis
To examine the bottlenecks in our ToL agent, we replace the local and global regions generated from the Hierarchical Layout Tree with human annotations, considered as the ground truth (GT). We also assess human performance by having human annotators write both content and layout descriptions manually. Due to the high cost of human efforts, we conduct experiments on a 20% random sample of the ScreenPR benchmark. As shown in Table 4, substituting either GT local regions or GT global regions significantly close the gap toward human performance in the ToL agent’s accuracy of content descriptions. However, for layout descriptions, the gap remains, indicating that the MLLM we leverage is the bottleneck in this aspect. This finding aligns with the conclusion from Fan et al. (2024), which highlights that current MLLMs face significant challenges in understanding the layout of multipanel images.
5.3 Broader Application: Verification of Mobile Navigation Agent Actions
In this section, we demonstrate how our ToL agent helps identify incorrect actions for a mobile navigation agent, MagicWonder, on MagicWand platformDing et al. . Driven by a specific task goal, such as "connect to Wifi in settings", MagicWonder can generate actions to interact with a live mobile screen and save the detailed information for each step along its execution trajectory. More details are provided in Appendix D.1. To incoorporate our ToL agent with the verification purpose, we design a simple pipeline. As shown in Figure 5, the layout and content descriptions are generated by our ToL agent for each action point and sent to the GPT-4-Turbo model. The GPT-4-Turbo model, based on the task goal, instructions, and descriptions from ToL agent, decides whether the last action from the MagicWonder is correct or not.
Data sampling
We select 52 mobile agent trajectories with specific task goals from instruction-guided executions on MagicWand platformDing et al. , with more details in Appendix D.2. Each trajectory consists of multiple execution steps triggered by distinct actions, resulting in a total action number of 209. We hire two students to annotate incorrect actions in the chosen trajectories to get the ground truth labels.
Result
We perform 10 end-to-end evaluations and present the mean value and standard deviation in Table 5. Despite minor fluctuations, while 92% correct actions can be stably recognized, about 38% incorrect actions can be identified. This reveals that incorporating the description from the ToL agent can significantly and robustly improve mobile navigation agents’ capability. Our ToL agent can also identify particularly tricky challenges faced by mobile navigation agents during execution. One notable issue is the "execution loop," where agents repeatedly take actions on the same or similar regions without making progress. Our results show that for all 20 incorrect actions fall into this category, 44% ± 0.04 can be successfully detected.
Predicted as TP/FP Rate
Correct Incorrect
Correct Action 100.2$\pm$2.0 8.8 $\pm$2.0 92% $\pm$0.02
Incorrect Action 61.9 $\pm$2.9 36.1 $\pm$2.9 38% $\pm$0.03
Table 5: Results of leveraging the ToL agent to verify actions from a mobile navigation agent. "Predicted as" column shows the number of predicted actions. True positive and false positive rate are shown on the right most column calculated from the data on the left
6 Conclusion and Discussion
In this work, we propose a novel Tree-of-Lens (ToL) grounding method and build a ToL agent for the Screen Point-and-Read (ScreenPR) task, an important screen referring task. Our ToL agent outputs layout-aware descriptions for pointed regions on screenshots. With our newly proposed Screen Point-and-Read (ScreenPR) benchmark, we show that our ToL agent outperforms all other baselines in terms of content and layout description accuracy, demonstrating great potential to improve accessibility of digital devices for visually impaired users. We also apply the ToL agent to identify incorrect actions in the execution trajectories of a mobile navigation agent, achieving promising results.
7 Limitations
In this work, we propose a novel Tree-of-Lens (ToL) agent for the Screen Point-and-Read (ScreenPR) task. Leveraging GPT4-o to generate content and layout descriptions allows us to set up the pipeline quickly and validate ToL’s effectiveness in a short time period. However, the accumulated server delay in requesting GPT4-o services has occupied more than 80% of end-to-end evaluation, not to mention the high expense accompanying, which is a non-trivial issue for real-world applications. Thus, a more efficient and locally hosted model has to be given out in place of GPT-4o. Moreover, since our ToL agent leverages MLLM to generate descriptions, there are potential risks that the generated descriptions from the MLLM could contain harmful contents. Therefore more post-processing strategy might be needed. Last but not least, another critical challenge is balancing application efficiency with improved semantic representation, thereby enabling on-device support and further expanding ToL’s application areas. Our current work is limited for research purpose and more application-wise exploration is pending.
Appendix A Example outputs of the ToL Agent
For each GUI domain, we show one example of the output of our ToL agent along with the lenses generated. We do not deliberately cherry pick the best case, but show examples with both strong and weak performances. An example for the operation system domain is shown in Figure 6. An example for the mobile domain is shown in Figure 7. An example for the web domain is shown in Figure 8.
Figure 6: Example of the generated description for an operation system screenshot from our ToL agent with the two lenses generated. The global region generated is not perfect as the bounding box is a little bigger than the application window on the top right part.
Figure 7: Example of the generated description for a mobile screenshot from our ToL agent with the two lenses generated.
Figure 8: Example of the generated description for a web screenshot from our ToL agent with the two lenses generated. The description output is not perfect regarding the layout information as the "product information area" indicated by the global region is not only "on the right half of the complete screen" but the middle and lower areas of the complete screen.
Appendix B Examples of the Cycle Consistency Evaluation
We select one result of our cycle consistency evaluation for the content description and show it in Figure 9. The left four images come from the cropped areas from different GUI screens, with only one of which matches the description in the prompt on the right side. As shown, GPT4-o can successfully choose the correct option in the "Answer" field and give its reasoning in the "Analysis" field.
Figure 9: One example of our cycle consistency evaluation for content description. A predicted description is shown on the right and four different cropped screen regions were sent as four separate images to GPT-4o. GPT-4o choose the matched region or answer with "unknown".
For cycle consistency evaluation regarding the layout description, the question contains the generated layout descriptions from models. We let GPT-4o choose one option from predefined 9 relative positions together with its analysis, referring to the layout evaluation output in Figure 11. The ground truth is provided with script based on the human annotated local regions for the target point and reference point. The ground truth is also checked by human.
Figure 10: Training Loss of ResNet-50 backbone with epoch 90.
Figure 11: One example of our cycle consistency evaluation for layout descriptions. The layout descriptions for target point and reference point are provided in prompt and GPT-4o is asked to decide their spatial relationship by choosing one option from the given 9 options.
{coloredquote}
Given the following information in a mobile navigation task:
Historical action and region description: {{haction}}
Task Goal: {{goal}}
Current region: {{region}}
The agent now is going to interact with the "Current region" with the action: {{action}}. Should the agent proceed?
Note: The agent should not proceed if the "Current region" is repeated too often in "Historical action and region description".
Note: The agent may proceed if the "Current region" aligns with "Task Goal".
Please provide your answer in the following JSON format:
"Analysis": "…",
"Answer": "yes/no"
Figure 12: Prompt used by GPT-4o in Action Validation
{coloredquote}
You are a smart screen reader that outputs concise natural language to answer questions from users based on the area (box 1) pointed out by the user shown as a red dot on the screen. The red dot is inside the box 1 in the first image, at (x,y) = ($\{\{x\}\}$,$\{\{y\}\}$), where x and y are the normalized coordinates.
Note: box 1 is the box with label 1 and box 2 is the box with label 2, box 1 is located inside box 2
Note: the first image shows the box 1 from the view of box 2, and the second image shows the box 2 from the complete screen.
Note: if the user asks about the location, based on the layout, explain where box 1 is in box 2 and then explain where box 2 is in the overall screen.
Note: don’t mention box 1, box 2 or the red dot in the output.
User question: (1) what is this? (2) where it is located in the screen?
Your output should in format (1) … (2) …
Figure 13: Prompt used by ToL agent
Prompt used by LlaVA-NEXT and GPT-4o
For the baseline models, LlaVA-NEXT and GPT-4o share the same prompting template as shown in Figure 14.
{coloredquote}
You are a smart screen reader that outputs concise natural language to answer questions from users based on the area pointed by the user at (x,y) = ($\{\{x\}\}$,$\{\{y\}\}$), where x and y are the normalized coordinates of the image.
Note: if the user asks about the location, based on the layout, explain where the target area is located locally and then explain where it is in the overall screen.
Note: don’t mention the area directly in the output, instead, use "it", "this".
User question: (1) what is this? (2) where it is located in the screen?
Your output should in format (1) … (2) …
Figure 14: Prompt used by LlaVA-NEXT and GPT-4o
Prompt used by CogAgent
In order to drive CogAgent to follow instructions, a simplified prompting template is applied, as shown in Figure 15.
{coloredquote}
What is indicated by the point at (x,y) = ($\{\{x\}\}$,$\{\{y\}\}$), where x and y are the normalized coordinates of the image.
and where it is located in the screen?
Figure 15: Prompt used by CogAgent
Figure 16: Given the goal of "pausing historial track in Youtube App", mobile Agent on MagicWand plans to click on the screen. ToL Agent generates descriptions for the action planed. The description is sent as input for action verification reasoning, where GPT-4o model output the analysis with decision, which is shown on the bottom right. The GPT-4o concludes that the click action is correct.
Appendix C Text Prompt templates
We show the templates of text prompts we used for different MLLMs. We substitute the $\{\{x\}\}$,$\{\{y\}\}$ with the normalized coordinate of the input point.
Prompt used by ToL agent
For our ToL agent, we use the text prompt template shown in Figure 13 for GPT-4o.
Appendix D Mobile Navigation Action Verification
D.1 Example
Figure 16 illustrates one action verification case in our experiment: MagicWonder is asked to accomplish "pause historical track in Youtube App". Len 1 and Len 2 show the agent was attempting to click the "Setting" icon by giving the instruction "Tap Settings" at the second step along the execution trajectory. The first step in action verification is to get the response from the ToL Agent that describes Len 1 and Len 2 correctly with more detailed application contexts. Using this information, GPT-4o could correctly judge whether this action on the screen was in accordance with the given instruction.
D.2 Sample Strategy of Mobile Navigation Action Execution Trajectory
MagicWand simulator can launch mobile agents to take action on devices under the guidance of a series of instructions. The corresponding runtime GUI context, like GUI screenshots, actionable regions, instructions for each step, and action details, are saved into a self-descriptive JSON format. We deliberately choose the instructions from the WeWeb dataset Ding et al. having few overlapping with our ASHL dataset. After execution on MagicWand, we chose those trajectories with task completion rates less than 50% to evaluate whether ToL information was accurate enough. The selected trajectories cover 11 Android applications from 8 application domains, where the chosen applications include SYSTEM, Entertainment, Education, Food & Drinks, Communication, Shopping, News, Maps & Navigation. The chosen Android apps include Messenger, Google Maps, YouTube, Settings, Quizlet, FlipBoard, McDonalds, eBay, Google Chat, Here WeGo, and DoorDash.
D.3 Action verification Pipeline Details
As shown in Figure 5, a set of navigation actions, target points, and corresponding screenshots, are first recorded during the interaction between the mobile navigation agent and the mobile screen. Then, our ToL agent first processes each pair of target points and screenshots $(P_{i},S_{i})$, including those from the history path of navigation trajectory and the currently planned one $(i\in[0,t-1])$, where the $P_{t}-1$ is the target point that the navigation agent plans to interact with for the next step. As a result, the layout and content descriptions $\hat{D}^{c}_{i}$, $\hat{D}^{l}_{i}$ are generated. Next, we concatenate all the descriptions and send them to the GPT-4-Turbo model to determine whether agent actions follow the given task goal.
Target path selection
For each mobile screenshot saved in the execution trajectory, the DINO detection model extracts all global regions with confidential scoring larger than 0.15 and local regions with confidential scoring larger than 0.05, consistent with our baseline evaluation. For click action, ToL first finds the smallest local and global regions that contain the click point. Usually, DINO’s output aligns with regions that have clear pixel boundaries. To improve the accuracy of local regions with smaller sizes, we extend DINO’s output regions by 50 pixels. If no match is found, we further extend the boundaries horizontally, following general GUI design principles. Different from click action, input action involves a region with a clear-cut boundary. Accordingly, ToL selects the smallest local region with an IoU value greater than 0.4 with the input region and the corresponding global region with an IoU value greater than 0.1 with the input region. With the chosen local and global regions, ToL follows Chain-of-Lens prompting mentioned in section 3.4 to get region description.
Action verification
We prompt GPT-4o by using the current action name, instruction, and region description at the current step, together with historical action names and region descriptions, letting it judge whether the mobile agent can proceed. Only two general rules are mentioned in the prompt:
  1. 1.
    The agent should not proceed if the "Current region" is repeated too often in "Historical action and region description"
  2. 2.
    The agent may proceed if the "Current region" aligns with "Instruction"
GPT4o’s response includes a simple yes or no and a detailed reason.
Prompt used by GPT-4o in Action Verification
To evaluate the role of LoT in mobile agents, we only list historical action and region description, current region, and the instruction for the current execution step together with the format we expect. During runtime, the aforementioned information will be used to replace variables in the prompting template {{haction}}, {{instruction}}, {{region}}, {{action}} separately.
We briefly mention two high-level goals that the agent is trying to achieve and expect it to figure out details via ToL information. More fine-grained work could be introduced later to get even better performance of action validation, though not our main focus now.
Appendix E Human Evaluation
We leverage human evaluation to test different screen reading models on our ScreenPR benchmark. We leverage the Amazon Mechanical Turk [^4] https://www.mturk.com platform to find annotators to complete the survey that we designed for our human evaluation. We pay \$0.15 for each survey and the screenshot of the survey is shown in Figure 18
Appendix F GUI Region Detection Model
Training details
Using MMDetection framework Chen et al. (2019), we change the training into two-label classification according to our task setting. Considering our dataset volume, we load pre-trained weights yielding the best performance on COCO, instead of training from scratch. We train the model for 90 epochs with a MultiStepLR scheduler set by milestones per 30 epochs and $\gamma$ = 0.1, the loss had further reduced to 2.5803 with small fluctuation, as shown in Figure 10.
We also examine the Average Precision metrics with different IoU and Area on the validation set: they show a clear convergence trend after epoch 70, and the results on the validation set are shown in Table 6.
IoU Area Average
0.50:0.95 all 0.9410
0.50 all 0.9620
0.75 all 0.9470
0.50:0.95 small 0.702
0.50:0.95 medium 0.897
0.50:0.95 large 0.943
Table 6: Average precision on the validation set for GUI region detection model. According to the definition in the Microsoft COCO dataset, the small and medium areas in an image are set as less than 1024 (32*32), 9216 (96*96), and others are large areas in the MMDetection framework.
Inference Examples and Evaluation Results
After training, we utilize the GUI region detection model on unseen OS, Web, and mobile screenshots in our ScreenPR benchmark. We show four examples of the detection results in Figure 17. The identified local regions are marked by green rectangles, and the global regions are highlighted in red. Additionally, we manually annotate global regions for 20% of randomly selected data from ScreenPR benchmark. With the human-annotated regions for the target points, we calculate the accuracy of our trained GUI region detection model as shown in Figure 7. Trained on only mobile data, our model shows generalization ability to other unseen domains.
Domains: Mobile Web OS Avg.
Local Region 38.4% 39.4% 10.0% 29.3%
Global Region 45.6% 35.8% 51.0% 44.1%
Table 7: Accuracy of the generated local region and global regions from our trained GUI region detection model compared with human annotation with IoU@0.5. Since the local regions are small, it brings a great challenge.
i Web page
ii Windows GUI
iii Ubuntu GUI
iv Unseen Mobile GUI
Figure 17: Inference results on Web, OS, and unseen Mobile GUI
Except for some undetected ones, many local and global regions were successfully identified with promising confidential scores. We assume that GUI elements share some common design patterns learned by DINO and guarantee good performance for zero-shot inference on heterogeneous platforms.
Figure 18: Screenshot of the survey we used in the human evaluation based on our ScreenPR benchmark. Due to the limited page length, questions are not shown completely, but the rest of the questions follow the same format as the one shown.